OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (anonymous):

@EclipsedStar@katti

OpenStudy (anonymous):

@katti

OpenStudy (anonymous):

Graph it first.. :)

OpenStudy (anonymous):

for a does 0+7=7 does 2+7=9 does 3+7=11?

OpenStudy (anonymous):

idk i need the answer plz

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i appreciate it

OpenStudy (anonymous):

yall are so kind